OpenStudy (anonymous):

What is an equation of the line, in point-slope form, that passes through the given point and has the given slope? point: (8, –8); slope: 3

OpenStudy (anonymous):

y-y1 = m (x - x1)

OpenStudy (anonymous):

(x1,y1) = (8,-8) and m=3

OpenStudy (anonymous):

y – 8 = 3 (x + 8) y + 8 = 3 (x + 8) y – 8 = 3 (x – 8) y + 8 = 3 (x – 8) which answer?

OpenStudy (anonymous):

last one